OpenStudy (anonymous):

Using an upper-case "C" for any arbitrary constants, find the general indefinite integral : integral |3 3root(x) dx|

OpenStudy (anonymous):

1. Take the 3 out of the integral. Always take constants out of the integral. 2. Assign u, which is x since it's derivative is 1dx, which is in the integral. 3. Take the antiderivative by adding 1 +1/3, which is 4/3. Take the 4/3 and put it in the denominator. The root of 3 = 1/3, which x is raised to. 4. The integral is gone, divide using keep change flip, and multiply across. 5. Your answer should be 9x^(4/3)/3 + C If that wasn't clear, let me know.

OpenStudy (anonymous):

The answer should have a denominator of 4, not 3.

OpenStudy (anonymous):

whats the denominator?

zepdrix (zepdrix):

Hmm this one might be a little bit trickier actually, due to the absolute bars. I'm assuming that's what those are, yes?

OpenStudy (anonymous):

zepdrix is right, is that an absolute value? I thought it was just to separate everything.

OpenStudy (anonymous):



OpenStudy (anonymous):

thats what the question looks like?

OpenStudy (anonymous):

ok, so the denominator is also 4/3.

OpenStudy (anonymous):

what is a denominator?

OpenStudy (anonymous):

do i get the antiderivative to work out the ans?

zepdrix (zepdrix):

Oh oh it's not an absolute :) That's a relief lol

OpenStudy (anonymous):

how do i work it out if its an absolute?

OpenStudy (anonymous):

When you have a fraction, there are two parts: the numerator and the denominator. The numerator is on top and the denominator is on bottom. The antiderivative is the same as the integral. Yes, take the antiderivative to get the answer.

OpenStudy (anonymous):

so the denominator is x? and the numerator is 3?

OpenStudy (anonymous):

how would i go about getting the antiderivative of this?

zepdrix (zepdrix):

\[\huge \int\limits3\; \sqrt[3]{x}\;dx\]This is what the problem looks like? :o

OpenStudy (anonymous):

ya :(

OpenStudy (anonymous):

I can't make the pictures so it's hard to explain, zeprdrix can so it's easier. After zepdrix helps you, there is a site called khanacademy.org that you can use. It's really good.

zepdrix (zepdrix):

I'll just restate some of Fringes steps with the math tool, maybe it will help :) Pull the constant outside of the integral, it won't affect the integration process.\[\large 3\int\limits \sqrt[3]{x}\;dx\]Rewrite the cube root, using a `rational expression`.\[\large 3\int\limits\limits x^{1/3}\;dx\] From here, we just apply the `Power Rule for Integration`. \[\large \color{royalblue}{\int\limits x^{n}dx\quad = \quad \frac{x^{n+1}}{n+1}}\]We increase the power by 1, then divide by the `new` power.

zepdrix (zepdrix):

\[\large 3\left(\frac{x^{(1/3+1)}}{\dfrac{1}{3}+1}\right)\]

zepdrix (zepdrix):

See how I put a +1 on the 1/3? It's always a little messy looking with all these fractions :C

OpenStudy (anonymous):

ya why did you put a +1?

zepdrix (zepdrix):

Remember how to take a derivative? It's 2 steps ~Multiply by the `old` power. ~Decrease the power by 1. With integration we're doing the opposite, and in the opposite order. ~Raise the power by 1. ~Divide by the `new` power.

zepdrix (zepdrix):

It might be easier to grasp the concept if you see an easy example,

OpenStudy (anonymous):

we do the antiderivative yes?

OpenStudy (anonymous):

so how would i go about finding the intergral with upper case C?

zepdrix (zepdrix):

Anti derivative is `Integration`. Yes :)

zepdrix (zepdrix):

That has been explained to you several times now :) lol Try to follow the steps posted above :O

OpenStudy (anonymous):

i dont understand it

OpenStudy (anonymous):

i think i figured it out!

zepdrix (zepdrix):

Here's an example.\[\large (x^3)'\]If we wanted to take the derivative of this, ~We wring the 3 (the power) down as multiplication. ~Then we -1 the power. \[\large 3x^{3-1}=3x^2\]______________________________________________ Now let's find the anti derivative of \(3x^2\).\[\large \int\limits 3x^2\;dx\]When we anti differentiate, the `big swirly S bar` and `dx` will disappear as part of this process. ~We'll +1 the power, ~Then we divide by the power. \[\large \frac{3x^{2+1}}{2+1} \quad = \quad \frac{3x^3}{3} \quad = x^3\]

zepdrix (zepdrix):

Seems like you're having some trouble with the idea of anti-derivatives. Just realize that it's the opposite process as taking a Derivative. See how we got back what we started with in this example?

OpenStudy (anonymous):